the length of a rectangle is three times its width. the perimeter is 24 ft. find the dimensions of the rectangle
Answer by deborabr(181)

width = x
length = 3x
perimeter is 2x + 2(3x)= 24
2x+6x=24
8x=24
x=24/8
x=3
width = 3 ft
length = 3*3 = 9 ft